OpenStudy (anonymous):

Which expression is rational? 6.27316543222718… square root of 2 square root of 14. square root of 49

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: the first number is an number with an infinity of decimal figures, so it is an irrational number, furthermore, we have 49=7^2, so: \[\Large \sqrt {49} = 7 = \frac{7}{1}\]

OpenStudy (michele_laino):

hint: a rational number R, is a number which can be written as a ratio between two integers, say m, and n, and such that n is different from zero, namely: \[\Large R = \frac{m}{n},\quad m,n \in \mathbb{Z},\;n \ne 0\]

OpenStudy (michele_laino):

\( \Large \sqrt 2 ,\quad \sqrt {14} \) are not expressible like a ratio above

OpenStudy (anonymous):

but they both can be rational

OpenStudy (michele_laino):

whereas: \[\Large \sqrt {49} = 7 = \frac{7}{1}\] is expressible as a ratio between 7 and 1

OpenStudy (michele_laino):

no, \( \large \sqrt 2 ,\quad \sqrt {14} \) are both irrational

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

oh i see they cant be expressed as a fraction

OpenStudy (michele_laino):

yes! So what is the right option?

OpenStudy (anonymous):

\[\sqrt{49}\]

OpenStudy (michele_laino):

that's right!